### v1

# A comprehensive review of assistive technologies for children with dyslexia

Abstract.
Dyslexia is a learning disability that primarily disrupts one’s ability to read, write and spell. About 15-20% of the world population is affected by dyslexia which signals it is a concerning subject. A systematic literature review was conducted on this research to evaluate trending interventions, common accessibility features, and potential barriers to existing assistive devices and to identify any research gaps. Papers published over period (2015-2023) of 8 years were reviewed and the data was meticulously analyzed. This research shows that technological-based interventions are leading interventions, especially with the use of mobile apps and augmentative realities. More innovative technologies like virtual reality, NLP, haptic, and tangible user interfaces are emerging to provide unique solutions addressing the user’s needs. Non-computing devices are less effective compared to modern digital solutions however, they provide a promising alternative.
Dyslexia, Children, Assistive Technology, Learning Disability
copyright: noneconference: ; ; copyright: none
1. Introduction
Dyslexia is a neurological disorder that primarily hinders one’s learning ability despite adequate intelligence, motivation, and educational expertise. It mainly affects one’s ability to read, write and spell [19]. According to international dyslexia association, dyslexia affects 15% to 20% of the world population[14], which means that around 1 in every 5 individuals has some form of dyslexia. Dyslexia is often misunderstood as a physical condition which in fact a myth. The particular region of human brain processes the smallest unit of sound called phonemes which is a fundamental unit for identifying and decoding words. The phonemes are automatically, segmented and mapped to particular words for normal individuals however, dyslexic children struggle with this ability due to a deficit in their phonological processing capability to break down written words into phonemes. This inefficiency hinders their reading fluency and word identification despite their intact critical thinking and reasoning skills. [36]. Children with dyslexia typically display several key attributes, including : (a) Difficulty in comprehending multiple instructions. (b)Difficulty paying attention, sitting still, listening to stories. (c) Difficulty with sequencing e.g. colored beads, classroom routines. (d) Difficulty learning nursery rhymes. Hence, Dyslexic children often avoids activities that involves reading[4]. There are no medications or proven therapies for dyslexia but special instruction can help people with dyslexia improve their literacy skills. The use of multi sensory technique, helps children learn through more than one sense is considered as an effective learning strategy for dyslexia, this technique demands visual, auditory, and kinesthetic activity [33, 32]. Various ranges of assistive technologies are prevalent to assists children with dyslexia. These include serious games, mobile apps, virtual reality, augmented reality, haptic devices, tactile tools, and non-computing devices. Apart from technological assistive tools which are more dominant non computing strategies also exists however, this approach often provides ordinary teaching and learning supports which lags engagement and is considered less optimal. The rise in technological device and literacy has resulted growth of technology in education. Technological advancement has open door to numerous possibilities on creating software products which are more interactive which can be tailored to individual needs.
This article provides the key findings on various functional elements that multiple technological and non computing device offers, with the key focus on multi-sensory methodology for categorization. The primary focus of current systematic literature survey includes identifying:
  • RQ1: What are the current interventions that enhance the learning experience for children with dyslexia, and what are their key features?
  • RQ2: What are the common accessibility features integrated into the existing interventions for children with dyslexia?
  • RQ3: What gaps exist in current assistive technology research for children with dyslexia, and what opportunities exist for innovation?
2. Background
Dyslexia is a lifelong learning disorder that primarily affects reading and spelling. Since, it affects reading, some people assumes dyslexia a problem with vision. However it is a problem with language processing. Reading starts with being able to recognize and work with the sounds of spoken language. This skill is called phonological awareness. Some of the common challenges faced by children with dyslexia include:
  • Trouble with phonics: Dyslexic children often don’t grasp phonics, which is a key skill for reading.
  • Slow to learn the alphabet: They may struggle to remember the order of letters or associate letters with their sounds.
  • Difficulty with sequencing: Tasks that require putting things in order, such as days of the week or steps in a process, are challenging.
  • Trouble decoding words: They find it hard to break words into sounds (phonemes) and blend them together.
  • Skipping or misreading words: They might miss small words like ”the” or confuse similar-looking words, such as ”was” and ”saw.”
  • Poor reading comprehension: Since reading takes so much effort, they might not understand what they’ve read, even if they can say the words.
  • Poor working memory: They have difficulty holding and manipulating information in their minds, such as remembering a phone number long enough to write it down.
  • Easily distracted: Because reading and writing are so challenging, they may lose focus or avoid tasks they find difficult.
  • Disorganized writing: Their writing may seem jumbled, with missing words, incorrect grammar, or random capitalization because they focus so much on spelling.
Multi-sensory activities are based on whole-brain learning, which is the belief that the best way to teach concepts is by involving multiple areas in the brain[25]. Adding auditory or visual components to reading assignments, like illustrations or online activities, helps students develop stronger literacy skills[23]. The benefits of multi-sensory learning have been verified by contemporary research in cognitive science [41]. Since, this model has been supported by multiple researchers one of my research goals in this literature survey was to evaluate existing technology with respect to multi-sensory approaches (Visual, Auditory, Kinesthetic, Tactile) in system design which fosters learning more engaging and interactive. The following sections contains key definitions that further elaborates the concept.
2.1. Visual
A visual can be anything to look at such as a picture, diagram, or piece of film that is used to make something more appealing or easier to understands. In the human brain, dyslexia primarily affects the visual processing areas, particularly impacting the ability to rapidly identify and sequence letters on a page, often due to disruptions in the ”magnocellular pathway” which controls visual attention and eye fixation, causing difficulties with accurately perceiving the visual details of words, leading to struggles with reading fluency[28].
2.2. Auditory
Audio is sound that is within the acoustic range of human hearing [37]. It is the powerful mechanism that can provide much detail information and insights for children with dyslexia. Children with dyslexia can learn and practice wide range of phonological sounds, distinguish similar sounds, sequencing sounds enhancing their learning abilities .
2.3. Kinesthetic
The term kinesthetic refers to the body movement that a learner uses to interact with their environment. The human kinesthetic system has the ability of processing, integrating, and remembering information, and the system is based on the immediate perception of motion impulses [35].
2.4. Tactile
The most complex and oldest-formed sensory system is the tactile sensation, which has the ability to translate information from the outside physical world into interior feelings [42].
3. Methodology
A systematic Literature was conducted by clearly defining inclusion and exclusion criteria, specifying database sources and developing a detailed search strategy with predefined keywords. To identify the recent technological advancement publications from 2015-2023 were selected for study.
3.1. Database source
The literature review was conducted through academic search engine which includes ACM, IEE and Google Scholar. The collection of journal articles from these engines allowed for a robust search strategy, capturing the most relevant papers from multiple reputable sources within the field of study.
3.2. Keyword search
The search for literature relevant to dyslexia and children was conducted using a structured boolean search strategy. The search string was divided into three key category, each representing a core concept: (1) the primary topic of interest, ”dyslexia,” (2) children or learning methods for children with dyslexia, and (3) tools or technologies utilized in these interventions. These groups were connected using the AND operator to retrieve studies that addressed all three aspects simultaneously. The following represents search string: (”dyslexia”) AND (”children” OR ”technology”) AND (”virtual reality” OR ”apps” OR ”mobile application” OR ”tactile” OR ”kinesthetic” OR ”auditory” OR ”games” OR ”haptic” OR ”educational games”).
3.3. Inclusion criteria
A set of criteria were defined to include research article for comprehensive review which are explained in table below.
Table 1. Inclusion Criteria
S.N Inclusion Criteria
1 Articles must be published between 2015-2024.
2 Articles must be written in English.
3 Publications must focus on children with dyslexia in the title or abstract.
4 Any visual, auditory, kinesthetic, or tactile approach used to facilitate the learning of pupils with dyslexia in any aspect.
3.4. Exclusion criteria
Exclusion criteria excludes the irrelevant journals that were out of scope to my study which were filtered out on the basis of following set of rules.
Table 2. Exclusion Criteria
S.N Criteria
1 Articles presenting only an abstract without the full document.
2 Research focus with poorly described abstract.
3 Articles related to dyslexia but not relevant to children or learning populations.
4 Articles focused on children with comorbid conditions.
5 Age group greater than 18.
3.5. Study selection
In order to ensure the findings, directly address the research questions, inclusion and exclusion criteria were introduced. This helps me select relevant publications within the field, and it acts as the initial step in my literature review. The search string was used as a query into popular libraries including IEEE, Google Scholar and ACM. Google Scholar returned 2430 results, ACM returned 515 and lastly, IEE resulted in 181 relevant results. After filtering the articles by year of publication between 1025 to 2024, Google Scholar returned 831, ACM library returned 312 and IEEE returned 96 results which were further processed for evaluation. The final 29 papers were selected after applying all sets of defined rules from inclusion and exclusion criteria. The detailed flow diagram is presented below.
Figure 1. Selection flow diagram
4. Emerging Assistive Technologies
A systematic literature review on assistive technologies for children with dyslexia was conducted, which resulted in assistive technologies being categorized into two primary categories: Technological and non-computational devices. Technological assistive devices encompass a wide range of technologies like augmented reality, virtual reality, mobile applications, computer applications, tangible user interface, and haptics whereas non-computing tools consist of various physical and structural methods of learning. Some non-computing techniques include learning with activities such as sky-writing, use of play-cards, use of magnifier, guide scale and many more. This diverse range of technologies provides multiple approaches to supporting children with dyslexia, offering both high-tech and classical solutions to address their unique learning needs.
5. Introduction to Technological Tools for Dyslexia
Technological tools are electronic, digital, or physical tools that can help people perform tasks [7]. There are wide range of technological practices designed for enhancing learning abilities for children with dyslexia. Many researchers conducted their research to resolve the learning disruption that children faced during pandemic due to COVID-19. A variety of technological devices are presented in the image below.
Figure 2. A:VR,B:Mobile app, C:Tangible user interface, D:AR
The following section will highlights key technical features that application provides along with their respective findings.
5.1. Mobile apps
A mobile application is a software application developed specifically for use with a mobile device such as a smartphone or tablet [31].In this literature review mobile application includes all the software applications particularly tailored for mobile and tablets including augmented reality.
5.1.1. Augmented reality
Augmented reality (AR) refers to the real-time integration of digital information into a user’s environment. AR technology overlays content onto the real world, enriching a user’s perception of reality rather than replacing it [29]. The researchers Hussain et al. on designing AR games for children utilized virtual objects to enhance the reading and writing skills of children with dyslexia. The application allows children to interact with 3D alphabets and solve word puzzles, and presents increasing level of difficulty to align with their cognitive development. This innovative approach can enhance active learning among children with dyslexia.
In this study Bhatti et al. leverages augmented reality to create multimedia learning application which uses AR markers or booklets to overlay 3D objects, to aid children in comprehending different educational elements such as shapes, letters, and words—tailored for children with dyslexia. Different educational practice exercises were implemented such as recognizing sounds, understanding different animals, and their educational course work from the specially designed educational booklet that supports AR for object identification and visualizations.
Another group of researchers Fei et al. from China had integrated augmented reality with computer vision technology to facilitate learning vocabulary and provided support for writing among children with dyslexia. This application provides children to engage in identifying real-world objects through augmented reality, where children scans real world object in their environment and the application produces audio cues for targeted object.CollectiAR infact, emphasizes the role of real-world engagement, encouraging children to interact with their environment to build vocabulary skills.
In this study Lazo-Amado and Andrade-Arenas created an augmented reality (AR) application to support children with dyslexia, focusing usability and accessibility. Researchers adopted user-centric design methodology to ensure the application meets the needs of its target audience. The application incorporates tasks such as word formation, letter identifications, word identification for augmented graphics, providing engaging and interactive learning experience. While the study reported positive usability feedback, it has limitation where the user study was conducted with parents of dyslexic children rather than children.
Table 3. Descriptive categorization of Augmented reality applications for children with dyslexia.
Category ARLexic AR Multimedia [5] CollectiAR Educadylexia
Research Gap Lack of engagement and interactivity in existing tools. Limited booklet design for AR. Existing word recognition and spelling training games were not able to teach students their nearby environments in real-time. Limited AR-based solutions for children with dyslexia in Latin America.
Technological Features Multiple Gaming Levels, word scramble game with timer, 3D alphabets and avatars, Rewards, Real-time alphabet identification for targeted physical object, graphics and animations, feedback support. AR-based multimedia learning using booklets as markers to overlay 3D objects, Audio cues with 3D graphics, Multiple exercises to identify letters, shapes, Numbers. Word hunt gameplay to find new words in physical environment, Audio cues to navigate target object, Provides animated feedback, Stores the captured images with words for future learning, Gameplay with spelling exercise,Tracks object and words collection for long term engagement support, Skip control Gaming Levels with 4 major activities, Completing Vowels, Selecting correct words based on image, Completing words based on image, Completing complex words without images
Personalization Adaptable levels. Not specified. Search control. Not Specified.
Targeted Device smartphone and tablets. smartphones and tablets. AR-enabled smartphones. Android-based smartphones and tablets.
Total Participants 21 children with age group of (7–14) years. Not specified but targeted for primary-level children. One child and two teachers in formative study. 80 (Parents)
Validation Test Paired sample t-test and Wilcoxon signed-rank test To Be Evaluated Formative study with qualitative feedback from teachers and children. Expert validation on Usability, Consistency, Functionality, Integration, resulting 98% acceptance rate
Findings Gamified activities such as word scramble effectively enhanced children engagement. This technology can be great help for enhanced cognitive learning via AR-based exercises. It has the potential to be a new type of supplementary vocabulary learning game for training children with dyslexia. Cognitive game improves the learning performance of children with dyslexia
Gaming Activities or Tasks Word games using 3D alphabets, puzzle games with increasing difficulty levels. Letter exercises with interactive visual and audio cues using AR booklets Vocabulary scavenger hunt using computer vision to identify and spell objects in surroundings. letter recognition, and comprehension tasks using AR overlays.
Software Used Unity 3D, Vuforia library for AR. Vuforia and Unity Used the computer vision model, SSD MobileNet V1. Figma, Tinkercad, MetaClassStudio.
Application tailored to augmented reality can further be categorized and evaluated based on multi-sensory modality, which implements a technique, (Visual, Auditory, Kinesthetic, and tactile). Since, multi sensory model is considered as one of the optimal model enhancing reading abilities. This high level categorization provides a quick reflection and implementation details to researchers on modalities used and targeted skills.The descriptive categorization of the application is presented on table-3.
Figure 3. Multi sensory modalities used on Augmented Reality.
The above table indicates that augmented reality application implements three different modalities which includes visual, auditory and kinesthetic (users movement) and represented to address mixed targeted skill sets which are crucial for enhancing the learning experience of children with dyslexia.
5.1.2. Mobile Applications
The researchers Rello et al. developed a mobile application designed to address most common spelling error and word formation for children with dyslexia. The natural language processing technique was used to process the collection of handwriting from children with dyslexia to identify the most relevant words on creating the game. Most common activities include adding, removing, and rearranging letters to create correct words. The application dynamically adjusts task difficulty based on individual performance, ensuring a tailored learning experience.
A group of researchers Vialatte et al. designed an application to enhance reading abilities through visual attentional training. Researchers developed symbol based visual search game where most alike symbols were presented with specific letters making identification challenging. This exercise was proven to be efficient in enhancing visual attention and reading accuracy among children with dyslexia.
The researchers Jaramillo-Alcázar et al. introduced a serious game tailored to support pattern recognition and enhancing spelling abilities.On this game design children interacts with the game play and their interaction are logged into database which are sent to server for further evaluation that provides therapist holistic view on analyzing the case.
The researchers Jun et al. designed game based learning system, focusing on enhancing spelling and letter identification skills. This game was designed distinguish between most alike letters (e.g., b/d and u/n) and improving overall reading abilities. Apart from this the application integrates dyslexia-friendly font and design principles to ensure accessibility and ease of use.
The researchers Bigueras et al. introduces game based learning with primary focus on enhancing letter and syllable recognition skills. This application provides story based learning which make it unique among all other applications.
The researchers Brennan et al. introduced mobile game with major focus on improving phonological awareness in children with dyslexia. This interactive game provides wide range of activities including exercises on developing phonological skills, such as blending of sound, identifying number of syllables, which is the foundational skills to excel reading abilities. It’s co-designed application hence, can provide more versatile learning experience.
The team of researchers Gupta et al. introduced a reading assistant designed to support children with dyslexia through real-time text interaction. Utilizing optical character recognition (OCR) and text-to-speech (TTS) technologies, the system provides customizable dyslexia-friendly features such as adjustable fonts, colors, and word spacing.
The following figure represents the multi-sensory modalities used on mobile applications, which highlights that mobile application heavily relies on visual and auditory cues. Furthermore mobile application addresses the demanding need on phonological awareness and reading practice among children with dyslexia.
Figure 4. Multi sensory modalities used on mobile apps.
Table 4. Descriptive categorization of mobile application for children with dyslexia
Categories Dyseggxia VisioCogLetters Puzzle Piece DysPReX LaroLexia Cosmic Sounds Augmenta11y
Research Gap Paper based exercises, which were limited in enhancing user experience. Limited visual-attention training for children with dyslexia. Limited cognitive games for children with dyslexia. Phonic-based intervention methods were ineffective, DysPReX provides solution with little emphasis on phonics. Children with dyslexia have different learning capability, existing tools are less efficient. Need for phonological awareness tools in dyslexia interventions. Limited real-time, personalized reading tools for dyslexic children.
Technological Feature Six types of word exercises for enhancing learning experience. Symbol-based visual search training game, Engaged children by challenging their symbol recognition and attentional skills. Multiple games to designed to practice word games., The game sends user’s game play interaction to server to recreate more robust system . Four variety game play with interactive interface, audio cues with major focuses on visual and auditory perception. Game based learning targeted to enhance reading performance in Filipino, specifically letter recognition and syllable reading using familiar Filipino words and sight-based learning techniques. Space games for phonological awareness. OCR-based reading assistant with customizable dyslexia-friendly features.
Personalization Adaptive exercise based on user’s performance. Not specified. Can select different game levels. customizable user settings. Tailored game elements and levels based on needs. Not Specified. Adjustable fonts, colors, word spacing, and TTS integration.
Targeted Device iOS and Android devices. Not specified. PC and mobile platforms. Mobile devices and tablets. Android devices. Not Specified Not Specified.
Cross-Platform Support Yes. Not specified. No, Android only. No, Android only. No, Android only. Not specified Yes, supports iOS and Android.
Total Participants 12 children with dyslexia. 21 dyslexic children. Targeted age group (5-10)years. Participant number not specified. 30 participants. 12 children, 5 teachers, and parents. 20 children with age group (9-12) years. Not Specified.
Validation Methods Not Specified ANOVA, Kendall Coeff. of Concordance, Wilcoxon tests. Not Specified. Usability and satisfaction scores from participants. Pre/post-test with paired t-test for reading performance. Wilcoxon signed-rank test for pre- and post-test results. Usability feedback and performance analysis.
Findings 5000 downloads on play-store,fun, engaging, and adaptive compared to traditional exercises. Improved visual search skills. Preliminary results suggests that accessibility features and customizable difficulty levels improved user experience, making the game suitable for children with dyslexia. Positive usability feedback on user interface, functionality, learnability, and satisfaction with scores above 4.0 out of 5. Significant improvement in reading performance. Improved phonological awareness skills and engagement. Increased engagement and improved reading experience.
Gaming Activities Word games including add a letter, remove a letter,change a letter, choose the correct word ending, cut into words. Identifying alike symbols with increasing difficulties. Identifying patterns on cards, Spell the visualized object. Spelling and letter differentiation games (e.g., b/d, u/n). Story based learning, points and rewards, multi level game play, feedback, and achievement system. Interactive Storyline Theme, Open-world exploration, Mission and Progressive Level, Rewards and acheivement, Feedback and Supportive Characters, Child-Centric Design. Scan and interact with textual materials in real time.
Software Used Not specified. Not specified. Unity game engine for mobile platforms, Firebase for real-time data storage and synchronization. Unity 3D. Not specified Not specified. OCR and TTS integrated into a custom app.
5.2. Tangible User Interface
A tangible user interface (TUI) is a user interface in which a person interacts with digital information through the physical environment. The initial name was Graspable User Interface, which is no longer used. The purpose of TUI development is to empower collaboration, learning, and design by giving physical forms to digital information, thus taking advantage of the human ability to grasp and manipulate physical objects and materials[15].
A group of researchers Jamali et al. introduced the use of tangible user interfaces to enhance learning engagement for children with dyslexia. Tangible objects like cards, toys were integrated with a mobile application, allowing children to interact and practice phonology, spelling, and reading. The engagement and fine motor skills involved in the tangible interaction promote a deeper understanding of sentence structure. The use of TUI with the blocks stimulates hands-on learning and facilitates the development of language skills, including grammar and syntax. The experiment with 30 children using this method resulted higher usability score i.e. 79.5The authors Antle et al. designed a tactile learning system integrating 3D printed tangible letters with dynamic color cues and tactile feedback to support letter recognition and sound association for dyslexic children.The printed letters, have mechanical keys that only particular letters fits on the device slots which helps children with dyslexia to identify mirror words like b and d. Researchers co-designed the activities with teachers to ensure alignment with educational needs and tailored the system to promote accurate letter orientation and sound decoding. While author had presented pilot project the real-world experiment are yet to be conducted however this system in in cooperates innovative approach to multi-sensory learning which helps in active learning among children with dyslexia.
Figure 5. Tangible user interface.
Table 5. Descriptive categorization of tangible User Interface for children with Dyslexia
Category DB3dT PhonoBlocks
Research Gap Limited tangible and multisensory teaching methodology for children with dyslexia. Not specified.
Technological Feature Tangible User Interface integrated with mobile application for enhanced learning engagement. Text to speech, color codes to identify letters sounds,helps in identifying mirrored letters like (p,q,d,b) to prevent spelling error and provides physical sensation to better understand alphabets.
Targeted Device Mobile tablets. Touch-screen laptop with an input platform and tangible letter blocks.
Targeted Skills Reading, spelling, pronunciation, and writing. Letter recognition, sound decoding, and rectifying mirror letter orientation.
Total Participants 30 children with dyslexia (15 in treatment group, 15 in control group). No participants yet, system co-designed with teachers for future validation.
Validation Methods System Usability Scale (SUS), Again-Again Table, observational notes, and performance checklist. Not specified.
Findings Improved engagement, learning performance, and usability score. Tangible letters with physical constraints may enable children to correctly use canonical orientation for letters, particularly for mirrored ones (p,q,d,b).
Gaming Activities Vowels, syllable, synonyms exercises, Interactive modules for phonology, spelling, and reading with tangible cards and toys. Activities involve arranging 3D printed letters into device slot to form words, with immediate feedback and dynamic color changes.
Software Used Not Specified. Arduino IDE for 3D letter processing and interaction.
5.3. Virtual Reality
Virtual Reality (VR) is a computer-generated environment with scenes and objects that appear to be real, making the user feel they are immersed in their surroundings. Researchers are exploring this noble domain to investigate its potential to enhance learning abilities in children with dyslexia.
A group of researchers Pedroli et al. introduced VR application for rehabilitation, particularly for children with dyslexia. The prime objective of this application was to enhance users engagement and adherence, to address the limitation traditional therapy posses. The application was tailored to make learning effective and enjoyable. Word identification, and retaining focus on children were major activities on this system. In summary significant improvements in word-reading test scores and homophonic writing signifying VR’s potential on rehabilitation an engaging alternative to traditional therapies.
The researchers Saunier et al. experimenting VR technology with Eye-Tracking (ET) and Brain-Computer Interface (BCI), to support children with dyslexia in therapy sessions. This prototype has not be clinically tested however author’s hypothesis on this technology is to monitor attention and cognitive load in real-time, thereby the system dynamically adapting exercises to the learner’s needs, enhancing personalization. This is a pilot study and the research is still on early phase however it can introduce novel method for combining neuroscience and VR to address the challenges faced by children with dyslexia.
A group of researchers Ullah et al., Am Faleel et al. has develop a customizable VR Learning Environment to each basic geometric shapes (cubes, cylinders, pyramids) to children with dyslexia. This application is assists children in in object identification and categorization tasks.Researchers had put forward multi-sensory VR realms which can bolster memory retention and comprehension through simultaneous sensory engagement. Dyslexic children can observe, listen to, and maneuver virtual entities, assisting them in grasping concepts more profoundly and cementing information more efficaciously.
The researchers Maresca et al. emphasizes multisensory learning on immersive VR environments. Researchers implemented user-centered designs to provide interactive reading and phonological tasks, fostering inclusion and motivation in learning processes. A significant difference in the word-reading test scores (p = 0.019) As well as in homophone writing (p = 0.034) was found.This qualitative analyses indicates higher engagement and motivation among users, presenting the effectiveness of multi-sensory VR interventions.
Virtual devices are used by children with dyslexia as a part of rehabilitation therapy. It is particularly used for object identification and adhering attention. The following table reflects the multi-sensory modalities used by virtual reality applications, where it encompasses all four elements.
Figure 6. Multi sensory modalities used on Virtual Reality devices.
Table 6. Descriptive categorization of virtual reality applications for children with dyslexia
Category Pedroli et al. Saunier et al. Ullah et al. Maresca et al.
Research Gap Traditional rehabilitation techniques lack user engagement and adherence due to repetitive tasks. Rehabilitation methods for dyslexia are repetitive and discourage therapy adherence. Limited multisensory techniques to address the challenges with dyslexia. Limited research on VR applications to enhance learning ability.
Technological Feature VR Rehabilitation System with prime focus on cognitive development. Multisensory exercises and biosensor technology,Hybrid VR interface combining eye tracking and brain computer interface. Customizable VR environment to teach object identification and categorization. Immersive and multisensory VR environments with prime focus on reading words.
Personalization Not Specified. Not Specified. Not Specified. Not Specified.
Targeted Device Not Specified. Not Specified. Not Specified. Not Specified.
Total Participants 10 children with dyslexia. Not Specified. Not Specified, Not Specified,designed for children aged 6-12. 28 children(14 control group and 14 experimental group).
Validation Methods Wechsler intelligence scale for children-IV. Pilot test. Qualitative analysis and feedback from participants. Mann–Whitney U-test.
Findings Significant decrease in the time of reading low frequency long words.Virtual Reality devices are efficient in rehabilitation for children with dyslexia on improving attentional skills Multisensory applications can be used for rehabilitation for children with dyslexia.Expected to improve therapy adherence and engagement. Improved engagement, object identification skills, and motivation. Virtual Reality in rehabilitation programs could be one assistive tool for children with dyslexia, classifying it as a promising treatment to enhance language skills, reduce disability, and promote psychological well-being.
Gaming Activities Posner cueing Task, Attentional Blink Task (identifying which letter is white), Reading words and not words(cat and lat). Gamified 3D exercises including to follow object to focus, finding Object in scene, reading exercise. Interactive categorization tasks using shapes and spatial relations. Exercises in which the children interacts with objects, Playful VR environment for motivation,homophone writing exercise.
Software Used Neurovirtual 3D. Unity. Not specified. Not specified.
5.4. Haptics
Haptic technology refers anything that can create an experience of touch by applying forces, vibrations, or motions to the user [38].
The researchers Park et al. introduced haptic technology that aids children with special abilities to improve fine motor control for their handwriting. Multilevel of haptic guidance were provided for various letter depending upon the complexity of alphabets. A group of 42 children participated in this research and the result showed that disturbance haptic guidance was the most effective for high complexity handwriting tasks (such as writing the letters “o” and “g”), partial haptic guidance was the most effective for medium complexity handwriting tasks (such as “t,” “r,” “s,” “e,” “n,” “a,” and “b”), and full haptic guidance was the most effective for low complexity letters (such as “i”). Haptic based technology provides real-time feedback with visual, kinesthetic, and touch sensation which can improve brain activity in learning. The Letters that were frequently used, confusing (b and d) were prioritized to evaluate haptic systems’ effectiveness in enhancing handwriting.
Similarly, Am Faleel et al. introduced Writely, a haptic force feedback system that uses Haptic force feedback for training non-dominant hand writing. Researchers had implemented guidance and anti-guidance system where the children who uses this system would provide a forced haptic feedback if they glides the letters trace. The system is on early stage of exploration, and can establish with a low cost solution for enhancing motor skills for special children.
Researchers Park et al. introduced falcon haptic device used in experiments. Children could grab the pencil-shaped end effector and the haptic device guides the participants’ handwriting to know how to write the task letter or draw task shape. 12 children with cognitive and fine motor delays and mild motor difficulty participated in the study. Haptic based learning guidance was resulted to improve motor functions. Although haptic device were not tailored to children with dyslexia, but all of the technology this technology can be a handy tool for children with dyslexia to enhance their learning abilities. The haptic devices presented all modalities, in enhancing writing abilities. The multi sensory model used on haptic device are shown in table below.
Figure 7. Multi sensory modalities used on haptic devices.
6. Introduction to Non-computing Tools for Dyslexia
Non computing assistive devices adopted traditional methodologies that existed prior to the advent of technological devices and although primitive, this methodology offers multi-sensory approach to guiding children for their effective learning. Researchers have implemented several non-computing strategies to enhance learning abilities of children with dyslexia, which is briefly described below.
The authors Ziadat purposes the VAKT modality which includes Visual, Auditory, kinesthetic and Tactile which provides multi sensory way of learning. The key purpose of the study was to examine how effective is this model in enhancing reading and reading comprehension among student with dyslexia. Several activities included during the experimental session such as word tracing, skywriting, and writing on the wrist. Apart form that word cards, color pens, crayon, mirror, sand, and whiteboard were used to make learning more elegant and interactive. Mirror was used a visual stimulating tool to look to his self when sounded out the phoneme, and simultaneously tapping his fingers to the thumb.39 students participated the survey which resulted statistically significant improvements in comprehension reading of the experimental group referred to the implemented multi-sensory approach (VAKT strategy).
A group of researchers Omar et al. from Malaysia conducted a comparison of visual aids for enhancing reading performance in children with dyslexia.Three specific tools were evaluated: a simple magnifier, which contains a magnifying lens to enlarge text; a visual tracking magnifier, which features a guide line along with a magnifying device that is effective in reducing glare while improving readability; and lastly, a typoscope, which focuses on a line of text by blocking out extra words and glare. 80 school children participated in this case study with critical finding that visual tracking magnifier(VTM) intervention yielded the greatest improvement after 12 weeks of intervention. Hence, author put forward an idea that a VTM can be used as part of a rehabilitation program for all children with dyslexia.
The authors Liu et al. introduced an structural approach to mitigate traditional method of teaching students with dyslexia. The practice of using the whole word to teach reading and writing for children with dyslexia were proven inefficient. The phonics-based program with special including structured, cumulative, multi-sensorial activities stimulate interest, attention as well as cultivate appropriate skills progressively. Several structural activities were designed to acheive the goal. Ranging activities from introduction of simple vowel, consonant letters to smaller phrases and sentence were introduced. Apart from this use of story-based Fitzroy books, laminated abc cards, a word list containing sequence of vowel, consonant alphabets, multi syllables exercise were introduced. Similarly the physical object or materials were used on this process which includes a set of bottle-caps representing each letter (26) in the alphabet in learning alphabets, tactile mat to practice skywriting, white board and whiteboard markers, exercise book, skipping ropes, hula hopes and many more. 28 total students participated on this experiment which resulted as the experiment had improved children progress in reading, spelling, writing and comprehension significantly determined by pre-test and post-test evaluation.
A group of researcher Ferraz et al. conducted an experiment to determine the effects of a phonological remediation on reading and writing program in children with dyslexia. Researchers evaluated phonological cal awareness, through multiple exercise like rapid naming, working memory, reading and writing of words and non words, thematic writing. Each phonological intervention consists of exercise involving phonemes and syllables, syllabic and phonemic manipulation, and rhyme and alliteration. The higher complex exercise includes the addition and subtraction of phonemes and syllables, syllabic and phonemic manipulation, oral comprehension etc.
The authors Heikkila and Knight explored inclusive music teaching strategies to enhance the learning abilities of elementary-aged children with developmental dyslexia. By incorporating rhythmic, melodic, and kinesthetic activities, the study aimed to improve phonological awareness, auditory discrimination, and motor coordination. Techniques included using body movements to represent pitches, clapping rhythms, and transitioning these skills to musical instruments like the keyboard. More descriptive categorization of non computing device is illustrated by the table below.
Table 7. Descriptive categorization of Non-computing devices and it’s qualities for dyslexic students
Non-computing Features Benefits
VAKT Strategy Combines Visual, Auditory, Kinesthetic, and Tactile learning (multisensory approach). Primary activities involved word tracing, skywriting, and writing on the wrist, playing with word cards, color pens, crayon, mirror, sand, and whiteboard for active learning. Improves oral reading and reading comprehension by engaging multiple senses.
Visual Tracking Magnifier (VTM) Magnifies text, provides a clear line guide, and reduces glare with filters. Enhances reading speed and accuracy, reduces visual strain, and helps with line focus.
Simple Magnifier Magnifies text without additional visual aids. Improves readability for children by enlarging text.
Typoscope Frames a specific line or area of text to reduce distractions. Focuses attention on a single line, improving reading comprehension and reducing cognitive load.
SMARTER Phonics Program Structured phonics-based reading program. Included multiple learning activites with the use of set of story-based Fitzroy books,laminated abc card, word list with syllables words, use of for learning alphabets, tactile mat for writing, skipping ropes and hula hoops for interactive learning Enhances decoding skills, phonemic awareness, and confidence in reading and writing, especially for rural and indigenous children.
Phonological Remediation Program Focused on phonological awareness, rapid naming, reading and writing of words and non words, thematic writing. Each phonological intervention consists of exercise involving phonemes, addition and subtraction of phoneme, syllables, syllabic and phone- mic manipulation, and rhyme and alliteration. The higher complex exercise includes the addition and subtraction of phonemes and syllables, syllabic and phonemic manipulation, oral comprehension Shows significant improvement in reading, writing, and memory retention through systematic intervention.
Music Therapy in Education Integrates melodic and rhythmic activities to engage students with activities including , auditory discrimination. Improves auditory processing, phonological skills, and temporal processing abilities, enhancing overall learning engagement.
7. Research Questions
7.1. RQ1:What are the current interventions that enhance the learning experience for children with dyslexia, and what are their key features?
The comprehensive literature survey resulted with finding that currently, there are two primary categories of interventions first being technological and second being non-computational. Technological intervention includes application ranging from low tech to high tech devices. These includes, audio books, mobile apps, augmented reality, virtual reality, haptic, tangible user interface, computer apps and natural language processing. Among all of these technological interventions mobile application was found to be most extensively researched domain considering the ease of development and usability among dyslexic children. Most application were focused on creating serious games or game based learning interventions however some application like tangible user interface were focused on providing more realistic way of learning in-cooperating all four (VAKT) multi-sensory parameters. Apart from these, modern immersive technologies like VR and AR systems, have utilized multi sensory approaches to improve letter recognition and phonological skills. Likewise, haptic-based tools have further enhance handwriting fluency and motor coordination through tactile feedback and adaptive guidance.
Non-computing interventions were more common on adopting traditional or prevailing methods. On this method the use of diverse activities such as use of assistive magnifiers, typoscope, play cards, wrist writing, musical games, sand writing etc were implemented that made learning interactive equally important in supporting dyslexic learners. Although classical, it adopts VAKT (Visual, Auditory, Kinesthetic, Tactile) strategy to reinforce memory and improve letter recognition through touch, sound, and movement. On this model structured word exercise were presented which resulted on positive outcome. Furthermore, non-computing intervention presents a challenge of requiring persistent guidance from teacher or caregiver, which makes it more dependent and less efficient comparing the use case of technological devices.
7.2. RQ2:What are the common accessibility features integrated into these interventions for children with dyslexia?
Taking both technological and non-computing devices into account most of the applications targeted the common challenge of dyslexia. This includes enhancing phonological awareness and letter-sound recognition. Immersive technologies like Virtual reality and augmented reality were focused on gamified activities that involved object’s and shape identification which were useful training for adhering attention. Non-computing devices provided learning platform, engaging with physical objects like books, colors, cards, whiteboard, it also contains structural exercise to bolster active learning. Different word games like syllabic identification, exercise related to identifying consonant and vowels sounds, spelling test, etc. were common in both technological and non-compositional devices. Many mobile applications were developed, with word games, taking account to the strategies of non-computing methodologies, thereby filling the gap it posses. Many technological devices presented audio cues as assistive features to enhance reading fluency and accuracy, similarly OCR and story based app were common in enhancing readability.
To sum up, gamified learning platform were most common features identified. Integration of rewards, and points to motivate users, real-time feedback were central idea almost every application implemented. Personalized features including customizable application, adjusting system settings such as font size, colors, and word spacing, making reading more comfortable and cross platform support were the next big thing that some application presented. Lastly, every application were implementing VAKT strategy in most possible way to enhance the learning abilities of children with dyslexia.
7.3. RQ3: What gaps exist in current assistive technology research for children with dyslexia, and what opportunities exist for innovation?
Current research on assistive technologies for children with dyslexia posses some gap. First being lack of long-term studies to evaluate the efficacy of this device, there is little to no evidence on whether these benefits persist over time. The existing research were conducted with small number participants, in some case even with their parents. This limits the generalizability of system across diverse population.
On other side, applications integrated with large language model or natural language processing were still limited. Their use case in children with dyslexia might result to a new finding which are yet to explore. Most mobile application were developed which were functional and limited to providing audio and visual cues, however no mobile application had designed an application in-cooperating haptic or vibratory feedback which could further bolster children’s learning, with activated sensory feedback. All this limits the widespread adoption and effectiveness of these technologies.
There were limited application which provided user’s personalization, hence personalized application could be the next domain to explore at. Despite these gaps, there are also some opportunities for innovation in this field to consider. First in-cooperating haptics feedback on mobile devices, secondly the use of LLM or NLP technology supporting to comprehend complex sentence in a more nuanced and simplified version. Lastly I further see potential in integration of AI on creating adaptive learning tools. AI could enable real-time analysis of a child’s performance and dynamically adjust tasks to tailor their skill level, ensuring personalized and effective learning experiences.
8. Discussion
In an exploration of the detailed literature review, I was fascinated by the use case of augmented reality in addressing the needs of dyslexic children. AR application are functional on most mobile or tablet devices, and popular being portable and affordable. The application provides 3D animated graphics, which make the learning environment more active, engaging and entertaining. AR are highly advance interm of functionality and popular for it’s mixed realities.
Although existing technologies for children with dyslexia uses smart and advance functionalities they possess some limitations. While, assistive technologies are expected to be both effective and affordable modern technological devices such as virtual reality, and haptic are expensive and complex, which might limit the reach of most children with dyslexia. Virtual device can further exacerbate eye strain with increased screen time. Likewise, growing children experience significant physiological changes, leading to fluctuating needs that can quickly render educational software obsolete or ineffective So, this barrier demands for more personalized applications that can solve the problem that growing children might face.
Similarly, applications that are tailored to learning disabilities possess varying font sizes, and bright background colors which can create a social stigma on use and can discourage them from using assistive technology which further can limit the use case of such devices.The findings from the systematic literature directs researchers or developers on creating more personalized, cross platform technologies in-cooperating multi-sensory modalities for recreating enhancing technologies.
9. Conclusion
The collective findings from the papers on assistive technologies for children with dyslexia highlights the importance of personalized applications to address unique learning challenges faced by children. Moreover, emerging technologies such as virtual reality, augmented reality, and haptic, tangible user interface has demonstrated a significant potential in enhancing and motivating children with learning disabilities. These technologies can provide a more realistic solution by incorporating all (VAKT), a multi-sensory model. Likewise, non-computing strategies, have proven to be efficient in reinforcing memory, for enhancing phonological awareness and can also be used as alternative method of learning where digital technologies are inaccessible.
Acknowledgements.
This Systematic Literature Review was completed as a part of my coursework. I extend my sincere gratitude to Professor Krishna Venkatasubramanian, for his guidance, support, and valuable insights throughout the semester.
Appendix A Appendix
List of abbreviations and acronyms:
S.N Abbreviation Description
1 3D Three Dimensional
2 SLR Systematic literature Review
3 AR Augmentative Reality
4 TUI Tangible User Interface
5 VR Virtual Reality
6 OCR Optical Character Recognition
7 VAKT Visual, Auditory, Kinesthetic, and Tactile
Table 8. List of Abbreviations and Descriptions