### v2

Abstract
Dyslexia is a neurological learning disability that primarily disrupts one’s ability to read, write and spell, affecting an estimated 15-20% of the global population. This high prevalence underscore the importance of developing effective interventions. This study presents a systematic literature review conducted between 2015 and 2024 to evaluate current trends in assistive technologies for children with dyslexia. This research shows that technological-based interventions are leading interventions, especially with the use of mobile apps and augmentative realities. More innovative technologies like virtual reality, NLP, haptic, and tangible user interfaces are emerging to provide unique solutions addressing the user’s needs. While non-computing devices are generally less effective in comparison to modern digital solutions, they provide a promising alternative in settings with limited access to technology.
keywords:
Dyslexia, Children Assistive Technology, Learning Disability, Technology Tools
1 Introduction
Dyslexia is a neurological disorder that primarily hinders one’s learning ability despite adequate intelligence, motivation, and educational expertise. It mainly affects one’s ability to read, write and spell[1]. According to international dyslexia association, dyslexia affects 15% to 20% of the world population[2], which means that around 1 in every 5 individuals has some form of dyslexia. Dyslexia is often misunderstood as a physical condition which in fact a myth. The particular region of human brain processes the smallest unit of sound called phonemes which is a fundamental unit for identifying and decoding words. The phonemes are automatically, segmented and mapped to particular words for normal individuals however, dyslexic children struggle with this ability due to a deficit in their phonological processing capability to break down written words into phonemes. This inefficiency hinders their reading fluency and word identification despite their intact critical thinking and reasoning skills. [3]. Children with dyslexia typically display several key attributes, including : (a) Difficulty in comprehending multiple instructions. (b) Difficulty paying attention, sitting still, listening to stories. (c) Difficulty with sequencing e.g. colored beads, classroom routines. (d) Difficulty learning nursery rhymes. Hence, Dyslexic children often avoids activities that involves reading[4]. There are no medications or proven therapies for dyslexia but special instruction can help people with dyslexia improve their literacy skills. The use of multi sensory technique, helps children learn through more than one sense is considered as an effective learning strategy for dyslexia, this technique demands visual, auditory, and kinesthetic activity [5, 6]. Various ranges of assistive technologies are prevalent to assists children with dyslexia. These include serious games, mobile apps, virtual reality, augmented reality, haptic devices, tactile tools, and non-computing devices. Apart from technological assistive tools which are more dominant non computing strategies also exists however, this approach often provides ordinary teaching and learning supports which lags engagement and is considered less optimal. The rise in technological device and literacy has resulted growth of technology in education. Technological advancement has open door to numerous possibilities on creating software products which are more interactive which can be tailored to individual needs.
This article provides the key findings on various functional elements that multiple technological and non computing device offers, with the key focus on multi-sensory methodology for categorization. The primary focus of current systematic literature survey includes identifying:
  • RQ1: What are the current interventions that enhance the learning experience for children with dyslexia, and what are their key features?
  • RQ2: What are the common accessibility features integrated into the existing interventions for children with dyslexia?
  • RQ3: What gaps exist in current assistive technology research for children with dyslexia, and what opportunities exist for innovation?
2 Background
Dyslexia is a lifelong learning disorder that primarily affects reading and spelling [7]. Since, it affects reading, some people assumes dyslexia a problem with vision. However it is a problem with language processing. Reading starts with being able to recognize and work with the sounds of spoken language. This skill is called phonological awareness. Some of the common challenges faced by children with dyslexia include:
  • Trouble with phonics: Dyslexic children often don’t grasp phonics, which is a key skill for reading.
  • Slow to learn the alphabet: They may struggle to remember the order of letters or associate letters with their sounds.
  • Difficulty with sequencing: Tasks that require putting things in order, such as days of the week or steps in a process, are challenging.
  • Trouble decoding words: They find it hard to break words into sounds (phonemes) and blend them together.
  • Skipping or misreading words: They might miss small words like ”the” or confuse similar-looking words, such as ”was” and ”saw.”
  • Poor reading comprehension: Since reading takes so much effort, they might not understand what they’ve read, even if they can say the words.
  • Poor working memory: They have difficulty holding and manipulating information in their minds, such as remembering a phone number long enough to write it down.
  • Easily distracted: Because reading and writing are so challenging, they may lose focus or avoid tasks they find difficult.
  • Disorganized writing: Their writing may seem jumbled, with missing words, incorrect grammar, or random capitalization because they focus so much on spelling.
Multi-sensory activities are based on whole-brain learning, which is the belief that the best way to teach concepts is by involving multiple areas in the brain[8]. Adding auditory or visual components to reading assignments, like illustrations or online activities, helps students develop stronger literacy skills[9]. The benefits of multi-sensory learning have been verified by contemporary research in cognitive science [10]. Since, this model has been supported by multiple researchers one of my research goals in this literature survey was to evaluate existing technology with respect to multi-sensory approaches (Visual, Auditory, Kinesthetic, Tactile) in system design which fosters learning more engaging and interactive. The following sections contains key definitions that further elaborates the concept.
2.1 Visual
A visual can be anything to look at such as a picture, diagram, or piece of film that is used to make something more appealing or easier to understands. In the human brain, dyslexia primarily affects the visual processing areas, particularly impacting the ability to rapidly identify and sequence letters on a page, often due to disruptions in the ”magnocellular pathway” which controls visual attention and eye fixation, causing difficulties with accurately perceiving the visual details of words, leading to struggles with reading fluency[11].
2.2 Auditory
Audio is sound that is within the acoustic range of human hearing [12]. It is the powerful mechanism that can provide much detail information and insights for children with dyslexia. Children with dyslexia can learn and practice wide range of phonological sounds, distinguish similar sounds, sequencing sounds enhancing their learning abilities .
2.3 Kinesthetic
The term kinesthetic refers to the body movement that a learner uses to interact with their environment. The human kinesthetic system has the ability of processing, integrating, and remembering information, and the system is based on the immediate perception of motion impulses [13].
2.4 Tactile
The most complex and oldest-formed sensory system is the tactile sensation, which has the ability to translate information from the outside physical world into interior feelings [14].
3 Methodology
A systematic Literature was conducted by clearly defining inclusion and exclusion criteria, specifying database sources and developing a detailed search strategy with predefined keywords. To identify the recent technological advancement publications from 2015-2024 were selected for study.
3.1 Database source
The literature review was conducted through academic search engine which includes ACM, IEEE and Google Scholar. The collection of journal articles from these engines allowed for a robust search strategy, capturing the most relevant papers from multiple reputable sources within the field of study.
3.2 Keyword search
The search for literature relevant to dyslexia and children was conducted using a structured boolean search strategy. The search string was divided into three key category, each representing a core concept: (1) the primary topic of interest, ”dyslexia,” (2) children or learning methods for children with dyslexia, and (3) tools or technologies utilized in these interventions. These groups were connected using the AND operator to retrieve studies that addressed all three aspects simultaneously. The following represents search string: (”dyslexia”) AND (”children” OR ”technology”) AND (”virtual reality” OR ”apps” OR ”mobile application” OR ”tactile” OR ”kinesthetic” OR ”auditory” OR ”games” OR ”haptic” OR ”educational games”).
3.3 Inclusion criteria
A set of criteria were defined to include research article for comprehensive review which are explained in Table 1
Table 1: Inclusion Criteria
S.N Inclusion Criteria
1 Articles must be published between 2015–2024.
2 Articles must be written in English.
3 Publications must focus on children with dyslexia in the title or abstract.
4 Any visual, auditory, kinesthetic, or tactile approach used to facilitate the learning of pupils with dyslexia in any aspect.
[…]
3.4 Exclusion criteria
Exclusion criteria excludes the irrelevant journals that were out of scope to my study which were filtered out on the basis of set of rules presented in Table 2.
Table 2: Exclusion Criteria
S.N Criteria
1 Articles presenting only an abstract without the full document.
2 Research focus with poorly described abstract.
3 Articles related to dyslexia but not relevant to children or learning populations.
4 Articles focused on children with comorbid conditions.
5 Age group greater than 18.
[…]
3.5 Study selection
In order to ensure the findings, directly address the research questions, inclusion and exclusion criteria were introduced. These criteria helped select relevant publications within the field and served as the initial step in the literature review. The search string was applied across major academic databases, including IEEE, Google Scholar, and ACM. Google Scholar initially returned 24200 results. From these, we selected top 1500 most recent publications. After screening for relevence, we excluded non-pertinent papers and finalized set of 1421 papers from this directory. In addition, ACM digital library returned 498 and IEEE Xplore resulted 151 relevent papers. Durign screening process 156 duplicates were identified across the databases, 32 papers were excluded for being non-English and 391 were excluded as they were not peer-reviewed. After filtering and applying the all above exclusion criteria the final 32 papers were selected after applying all sets of defined rules from inclusion and exclusion criteria. The detailed flow diagram is presented in Figure 1.
Figure 1: Article selection process based on inclusion and exclusion criteria. The flow diagram illustrates the screening of articles from three digital libraries (ACM, Google Scholar, and IEEE)
4 Emerging Assistive Technologies
A systematic literature review on assistive technologies for children with dyslexia was conducted, which resulted in assistive technologies being categorized into two primary categories: Technological and non-computational devices. Technological assistive devices encompass a wide range of technologies like augmented reality, virtual reality, mobile applications, computer applications, tangible user interface, and haptics whereas non-computing tools consist of various physical and structural methods of learning. Some non-computing techniques include learning with activities such as sky-writing, use of play-cards, use of magnifier, guide scale and many more. This diverse range of technologies provides multiple approaches to supporting children with dyslexia, offering both high-tech and classical solutions to address their unique learning needs.
5 Introduction to Technological Tools for Dyslexia
Technological tools are electronic, digital, or physical tools that can help people perform tasks [15]. There are wide range of technological practices designed for enhancing learning abilities for children with dyslexia. Many researchers conducted their research to resolve the learning disruption that children faced during pandemic due to COVID-19. A variety of technological devices are presented in Figure 2.
Figure 2: Examples of assistive technologies used in interventions for dyslexia: (A) Virtual Reality environment, (B) Mobile learning applications, (C) Tangible user interface with illuminated letter blocks, and (D) Augmented Reality learning using 3D visualizations.
The following section will highlights key technical features that application provides along with their respective findings.
5.1 Mobile-Assisted Learning Applications
Mobile-assisted learning applications are software applications designed to be used with mobile devices like smartphones or tablets [16]. In this literature review mobile-assisted learning application includes all the software applications particularly tailored for mobile and tablets including augmented reality.
5.1.1 Augmented reality (AR)
AR refers to the real-time integration of digital information into a user’s environment. AR technology overlays content onto the real world, enriching a user’s perception of reality rather than replacing it [17]. The researchers Hussain et al. [18] on designing AR games for children utilized virtual objects to enhance the reading and writing skills of children with dyslexia. The application allows children to interact with 3D alphabets and solve word puzzles, and presents increasing level of difficulty to align with their cognitive development. This innovative approach can enhance active learning among children with dyslexia.
In this study Bhatti et al. [19] leverages augmented reality to create multimedia learning application which uses AR markers or booklets to overlay 3D objects, to aid children in comprehending different educational elements such as shapes, letters, and words—tailored for children with dyslexia. Different educational practice exercises were implemented such as recognizing sounds, understanding different animals, and their educational course work from the specially designed educational booklet that supports AR for object identification and visualizations.
Another group of researchers Fei et al. [20] from China had integrated augmented reality with computer vision technology to facilitate learning vocabulary and provided support for writing among children with dyslexia. This application provides children to engage in identifying real-world objects through augmented reality, where children scans real world object in their environment and the application produces audio cues for targeted object. CollectiAR infact, emphasizes the role of real-world engagement, encouraging children to interact with their environment to build vocabulary skills.
In this study Amado and Arenas et al. [21] created an augmented reality (AR) application to support children with dyslexia, focusing usability and accessibility. Researchers adopted user-centric design methodology to ensure the application meets the needs of its target audience. The application incorporates tasks such as word formation, letter identifications, word identification for augmented graphics, providing engaging and interactive learning experience. While the study reported positive usability feedback, it has limitation where the user study was conducted with parents of dyslexic children rather than children.
Table 3: Descriptive categorization of Augmented Reality applications for children with dyslexia.
Category ARLexic AR Multimedia [19] CollectiAR Educadylexia
Research Gap Lack of engagement and interactivity in existing tools. Limited booklet design for AR. Existing word recognition and spelling training games were not able to teach students their nearby environments in real time. Limited AR-based solutions for children with dyslexia in Latin America.
Technological Features Multiple gaming levels; word-scramble with timer; 3D alphabets and avatars; rewards; real-time alphabet identification for targeted physical objects; graphics and animations; feedback support. AR-based multimedia learning using booklets as markers to overlay 3D objects; audio cues with 3D graphics; multiple exercises to identify letters, shapes, and numbers. Word-hunt gameplay to find new words in the physical environment; audio cues to navigate to target objects; animated feedback; stores captured images with words for future learning; gameplay with spelling exercise; tracks objects/words collection for long-term engagement; skip control. Gaming levels with four activities: completing vowels; selecting correct words based on image; completing words based on image; completing complex words without images.
Personalization Adaptable levels. Not specified. Search control. Not specified.
Targeted Device Smartphones and tablets. Smartphones and tablets. AR-enabled smartphones. Android-based smartphones and tablets.
Total Participants 21 children (ages 7–14). Not specified; primary-level children targeted. One child and two teachers (formative study). 80 (parents).
Validation Test Paired-sample t-test and Wilcoxon signed-rank test. To be evaluated. Formative study with qualitative feedback from teachers and children. Expert validation on usability, consistency, functionality, and integration; 98% acceptance rate.
Findings Gamified activities such as word scramble enhanced engagement. AR-based exercises can enhance cognitive learning. Potential supplementary vocabulary-learning game for training children with dyslexia. Cognitive game improves learning performance of children with dyslexia.
Gaming Activities or Tasks Word games using 3D alphabets; puzzle games with increasing difficulty. Letter exercises with interactive visual and audio cues using AR booklets. Vocabulary scavenger hunt using computer vision to identify and spell objects in surroundings. Letter recognition and comprehension tasks using AR overlays.
Software Used Unity 3D; Vuforia library for AR. Vuforia; Unity. Computer-vision model: SSD MobileNet V1. Figma; Tinkercad; MetaClassStudio.
[…]
Application tailored to augmented reality can further be categorized and evaluated based on multi-sensory modality, which implements a technique (visual, auditory, kinesthetic, and tactile). Since, multi sensory model is considered as one of the optimal model enhancing reading abilities. This high level categorization provides a quick reflection and implementation details to researchers on modalities used and targeted skills. The descriptive categorization of the application is presented on Table LABEL:tab:AR_comparison_dyslexia.
Figure 3: Multi sensory modalities used on Augmented Reality.
The Figure 3 indicates that augmented reality application implements three different modalities which includes visual, auditory and kinesthetic (users movement) and represented to address mixed targeted skill sets which are crucial for enhancing the learning experience of children with dyslexia.
5.1.2 Mobile Applications
The researchers Rello et al. [22] developed a mobile application designed to address most common spelling error and word formation for children with dyslexia. The natural language processing technique was used to process the collection of handwriting from children with dyslexia to identify the most relevant words on creating the game. Most common activities include adding, removing, and rearranging letters to create correct words. The application dynamically adjusts task difficulty based on individual performance, ensuring a tailored learning experience.
A group of researchers Vialatte et al. [23] designed an application to enhance reading abilities through visual attentional training. Researchers developed symbol based visual search game where most alike symbols were presented with specific letters making identification challenging. This exercise was proven to be efficient in enhancing visual attention and reading accuracy among children with dyslexia.
The researchers Jaramillo-Alc´azar et al. [24] introduced a serious game tailored to support pattern recognition and enhancing spelling abilities.On this game design children interacts with the game play and their interaction are logged into database which are sent to server for further evaluation that provides therapist holistic view on analyzing the case.
The researchers Jun et al. [25] designed game based learning system, focusing on enhancing spelling and letter identification skills. This game was designed distinguish between most alike letters (e.g., b/d and u/n) and improving overall reading abilities. Apart from this the application integrates dyslexia-friendly font and design principles to ensure accessibility and ease of use.
The researchers Bigueras et al. [26] introduces game based learning with primary focus on enhancing letter and syllable recognition skills. This application provides story based learning which make it unique among all other applications.
The researchers Brennan et al. [27] introduced mobile game with major focus on improving phonological awareness in children with dyslexia. This interactive game provides wide range of activities including exercises on developing phonological skills, such as blending of sound, identifying number of syllables, which is the foundational skills to excel reading abilities. It’s co-designed application hence, can provide more versatile learning experience.
The team of researchers Gupta et al. [28] introduced a reading assistant designed to support children with dyslexia through real-time text interaction. Utilizing optical character recognition (OCR) and text-to-speech (TTS) technologies, the system provides customizable dyslexia-friendly features such as adjustable fonts, colors, and word spacing.
The Figure 4 represents the multi-sensory modalities used on mobile applications, which highlights that mobile application heavily relies on visual and auditory cues. Furthermore mobile application addresses the demanding need on phonological awareness and reading practice among children with dyslexia. The descriptive categorization is presented in Table 4.
Figure 4: Multi sensory modalities used on mobile apps.
Table 4: Descriptive categorization of mobile applications for children with dyslexia.
Categories Dyseggxia VisioCogLetters Puzzle Piece DysPReX
Research Gap Paper-based exercises limited engagement. Limited visual-attention training for dyslexic children. Few cognitive games aimed at dyslexia. Phonics-based methods ineffective; DysPReX reduces phonics focus.
Technological Features Six types of word exercises. Symbol-based visual search training; challenges symbol recognition and attention. Multiple word games; gameplay interactions logged to server. Four gameplay types; interactive UI; audio cues; focus on visual and auditory perception.
Personalization Adaptive exercises based on performance. Not specified. Selectable levels. Customizable user settings.
Targeted Device iOS and Android. Not specified. PC and mobile. Mobile and tablets.
Cross-Platform Support Yes. Not specified. No (Android only). No (Android only).
Total Participants 12 children with dyslexia. 21 dyslexic children. Ages 5–10 (count not specified). 30 participants.
Validation Methods Not specified. ANOVA; Kendall’s W; Wilcoxon. Not specified. Usability and satisfaction scores.
Findings 5000+ downloads; engaging and adaptive vs. paper exercises. Improved visual search skills. Accessibility features + adjustable difficulty improved UX. Positive usability (UI, functionality, learnability, satisfaction >4/5>4/5).
Gaming Activities Add/remove/change letter; choose endings; cut into words. Identify alike symbols with increasing difficulty. Pattern recognition; spell visualized object. Spelling and letter-differentiation (e.g., b/d, u/n).
Software Used Not specified. Not specified. Unity + Firebase. Unity 3D.
[…]
Table 5: Descriptive categorization of mobile applications for children with dyslexia (continued).
Categories LaroLexia Cosmic Sounds Augmenta11y Cahyani et al. [29]
Research Gap Different learning capabilities; existing tools less efficient. Need for phonological-awareness tools in interventions. Limited real-time, personalized reading tools. Limited integration of user-centred UI with gamification and multi-sensory learning.
Technological Features Game-based reading (letter recognition, syllable reading, sight words) using familiar Filipino words. Space-themed games for phonological awareness. OCR-based reading assistant with customizable dyslexia-friendly settings. Mobile UI via design thinking; guided menus, learning paths, visual/audio/kinesthetic integration, gamified tasks, adaptive progress flow.
Personalization Tailored elements and levels. Not specified. Adjustable fonts, colours, spacing; TTS integration. Gradual content unlocking by progress.
Targeted Device Android. Not specified. Not specified. Android mobile/tablet (Figma prototype).
Cross-Platform Support No (Android only). Not specified. Yes (iOS and Android). Not specified.
Total Participants 12 children, 5 teachers, and parents. 20 children (ages 9–12). Not specified. 10 dyslexic students.
Validation Methods Pre/post paired t-test for reading performance. Wilcoxon signed-rank (pre/post). Usability feedback and performance analysis. System Usability Scale: 82.
Findings Significant reading improvement. Improved phonological awareness and engagement. Increased engagement and improved reading experience. Usable, inclusive; supports structured reading/counting practice.
Gaming Activities Story-based, multilevel; feedback; rewards; achievements. Interactive storyline; missions; progressive levels; rewards; supportive characters. Scan and interact with printed text in real time. Guessing alphabets/words; ranking modes; stepwise access to activities.
Software Used Not specified. Not specified. OCR + TTS custom app. Figma.
[…]
5.2 Tangible User Interface
A tangible user interface (TUI) is a user interface in which a person interacts with digital information through the physical environment. The initial name was Graspable User Interface, which is no longer used. The purpose of TUI development is to empower collaboration, learning, and design by giving physical forms to digital information, thus taking advantage of the human ability to grasp and manipulate physical objects and materials[30].
A group of researchers Jamali et al. [31] introduced the use of tangible user interfaces to enhance learning engagement for children with dyslexia. Tangible objects like cards, toys were integrated with a mobile application, allowing children to interact and practice phonology, spelling, and reading. The engagement and fine motor skills involved in the tangible interaction promote a deeper understanding of sentence structure. The use of TUI with the blocks stimulates hands-on learning and facilitates the development of language skills, including grammar and syntax. The experiment with 30 children using this method resulted higher usability score i.e. 79.5%.
The authors Antle et al. [32] designed a tactile learning system integrating 3D printed tangible letters with dynamic color cues and tactile feedback to support letter recognition and sound association for dyslexic children. The printed letters, have mechanical keys that only particular letters fits on the device slots which helps children with dyslexia to identify mirror words like ’b’ and ’d’. Researchers co-designed the activities with teachers to ensure alignment with educational needs and tailored the system to promote accurate letter orientation and sound decoding. While author had presented pilot project the real-world experiment are yet to be conducted however this system in in cooperates innovative approach to multi-sensory learning which helps in active learning among children with dyslexia.
TUI-based interventions are summarized in Figure 5, highlighting the multisensory modalities they employ (visual, auditory, kinesthetic, tactile) and the specific language skills they target and descriptive details in Table 6
Figure 5: Tangible user interface.
Table 6: Descriptive categorization of tangible user interfaces for children with dyslexia.
Category DB3dT PhonoBlocks
Research Gap Limited tangible and multisensory teaching methodology for children with dyslexia. Not specified.
Technological Features Tangible user interface integrated with a mobile application for enhanced learning engagement. Text-to-speech; colour-coded letter–sound mapping; support for mirrored letters (p, q, d, b) to reduce spelling errors; physical interaction to reinforce letter understanding.
Targeted Device Tablet devices. Touch-screen laptop with an input platform and tangible letter blocks.
Targeted Skills Reading, spelling, pronunciation, and writing. Letter recognition, sound decoding, and rectifying mirror-letter orientation.
Total Participants 30 children with dyslexia (15 treatment, 15 control). No participants yet; system co-designed with teachers for future validation.
Validation Methods System Usability Scale (SUS); Again–Again Table; observational notes; performance checklist. Not specified.
Findings Improved engagement, learning performance, and usability scores. Physical letter constraints may help children use canonical letter orientation, particularly for mirrored letters (p, q, d, b).
Gaming Activities Vowel, syllable, and synonym exercises; interactive modules for phonology, spelling, and reading using tangible cards and toys. Arrange 3D-printed letters into the device slot to form words, with immediate feedback and dynamic colour changes.
Software Used Not specified. Arduino IDE for 3D-letter processing and interaction.
[…]
5.3 Virtual Reality
Virtual Reality (VR) is a computer-generated environment with scenes and objects that appear to be real, making the user feel they are immersed in their surroundings. Researchers are exploring this noble domain to investigate its potential to enhance learning abilities in children with dyslexia.
A group of researchers Pedroli et al. [33] introduced VR application for rehabilitation, particularly for children with dyslexia. The prime objective of this application was to enhance users engagement and adherence, to address the limitation traditional therapy posses. The application was tailored to make learning effective and enjoyable. Word identification, and retaining focus on children were major activities on this system. In summary significant improvements in word-reading test scores and homophonic writing signifying VR’s potential on rehabilitation an engaging alternative to traditional therapies.
The researchers Saunier et al. [34] experimenting VR technology with Eye-Tracking (ET) and Brain-Computer Interface (BCI), to support children with dyslexia in therapy sessions. This prototype has not be clinically tested however author’s hypothesis on this technology is to monitor attention and cognitive load in real-time, thereby the system dynamically adapting exercises to the learner’s needs, enhancing personalization. This is a pilot study and the research is still on early phase however it can introduce novel method for combining neuroscience and VR to address the challenges faced by children with dyslexia.
A group of researchers [35, 36] has develop a customizable VR Learning Environment to each basic geometric shapes (cubes, cylinders, pyramids) to children with dyslexia. This application is assists children in in object identification and categorization tasks.Researchers had put forward multi-sensory VR realms which can bolster memory retention and comprehension through simultaneous sensory engagement. Dyslexic children can observe, listen to, and maneuver virtual entities, assisting them in grasping concepts more profoundly and cementing information more efficaciously.
The researchers Maresca et al. [37] emphasizes multisensory learning on immersive VR environments. Researchers implemented user-centered designs to provide interactive reading and phonological tasks, fostering inclusion and motivation in learning processes. A significant difference in the word-reading test scores (p = 0.019) as well as in homophone writing (p = 0.034) was found. This qualitative analyses indicates higher engagement and motivation among users, presenting the effectiveness of multi-sensory VR interventions.
Virtual devices are used by children with dyslexia as a part of rehabilitation therapy. It is particularly used for object identification and adhering attention. The Figure 6 reflects the multi-sensory modalities used by VR applications, where it encompasses all four elements with details in Table 7
Figure 6: Multi sensory modalities used on Virtual Reality devices.
Table 7: Descriptive categorization of virtual reality applications for children with dyslexia.
Category Pedroli et al. [33] Saunier et al. [34] Ullah et al. [35] Maresca et al. [37]
Research Gap Traditional rehabilitation techniques are repetitive, reducing engagement and adherence. Rehabilitation for dyslexia tends to be repetitive and discourages therapy adherence. Limited multisensory techniques addressing dyslexia-specific challenges. Limited research on VR to enhance learning ability.
Technological Features VR rehabilitation system with a focus on cognitive development. Multisensory exercises with biosensors; hybrid VR interface combining eye tracking and brain–computer interface. Customizable VR environment to teach object identification and categorization. Immersive, multisensory VR environments with a focus on reading words.
Personalization Not specified. Not specified. Not specified. Not specified.
Targeted Device Not specified. Not specified. Not specified. Not specified.
Total Participants 10 children with dyslexia. Not specified. Designed for children aged 6–12 (participant count not specified). 28 children (14 control, 14 experimental).
Validation Methods Wechsler Intelligence Scale for Children (WISC–IV). Pilot test. Qualitative analysis and participant feedback. Mann–Whitney U-test.
Findings Significant decrease in time to read low-frequency long words; VR devices improved attentional skills during rehabilitation. Multisensory VR can support rehabilitation for dyslexia; expected to improve adherence and engagement. Improved engagement, object-identification skills, and motivation. VR in rehabilitation shows promise for improving language skills, reducing disability, and promoting psychological well-being.
Gaming Activities Posner cueing task; attentional blink task (identify the white letter); reading words and non-words (e.g., cat/lat). Gamified 3D exercises: following/focusing on moving objects; object-finding; reading exercises. Interactive categorization using shapes and spatial relations. Object interaction in playful VR scenes; homophone writing exercise to maintain motivation.
Software Used NeuroVirtual 3D. Unity. Not specified. Not specified.
[…]
5.4 Haptics
Haptic technology refers anything that can create an experience of touch by applying forces, vibrations, or motions to the user [38].
The researchers Park et al. [39] introduced haptic technology that aids children with special abilities to improve fine motor control for their handwriting. Multilevel of haptic guidance were provided for various letter depending upon the complexity of alphabets. A group of 42 children participated in this research and the result showed that disturbance haptic guidance was the most effective for high complexity handwriting tasks (such as writing the letters “o” and “g”), partial haptic guidance was the most effective for medium complexity handwriting tasks (such as “t,” “r,” “s,” “e,” “n,” “a,” and “b”), and full haptic guidance was the most effective for low complexity letters (such as “i”). Haptic based technology provides real-time feedback with visual, kinesthetic, and touch sensation which can improve brain activity in learning. The Letters that were frequently used, confusing (b and d) were prioritized to evaluate haptic systems’ effectiveness in enhancing handwriting.
Similarly, Am Faleel et al. [36] introduced Writely, a haptic force feedback system that uses Haptic force feedback for training non-dominant hand writing. Researchers had implemented guidance and anti-guidance system where the children who uses this system would provide a forced haptic feedback if they glides the letters trace. The system is on early stage of exploration, and can establish with a low cost solution for enhancing motor skills for special children.
Researchers Park et al. [8] introduced falcon haptic device used in experiments. Children could grab the pencil-shaped end effector and the haptic device guides the participants’ handwriting to know how to write the task letter or draw task shape. 12 children with cognitive and fine motor delays and mild motor difficulty participated in the study. Haptic based learning guidance was resulted to improve motor functions. Although haptic device were not tailored to children with dyslexia, but all of the technology this technology can be a handy tool for children with dyslexia to enhance their learning abilities. The haptic devices presented all modalities, in enhancing writing abilities. The multi sensory model used on haptic device are shown in Figure 7.
Figure 7: Multi sensory modalities used on haptic devices.
6 Introduction to Non-computing Tools for Dyslexia
Non computing assistive devices adopted traditional methodologies that existed prior to the advent of technological devices and although primitive, this methodology offers multi-sensory approach to guiding children for their effective learning. Researchers have implemented several non-computing strategies to enhance learning abilities of children with dyslexia, which is briefly described below.
The authors Ziadat [40] purposes the VAKT modality which includes Visual, Auditory, kinesthetic and Tactile which provides multi sensory way of learning. The key purpose of the study was to examine how effective is this model in enhancing reading and reading comprehension among student with dyslexia. Several activities included during the experimental session such as word tracing, skywriting, and writing on the wrist. Apart form that word cards, color pens, crayon, mirror, sand, and whiteboard were used to make learning more elegant and interactive. Mirror was used a visual stimulating tool to look to his self when sounded out the phoneme, and simultaneously tapping his fingers to the thumb. Thirty-nine students participated in the survey, which revealed statistically significant improvements in comprehension reading for the experimental group that utilized the implemented multi-sensory approach (VAKT strategy).
A group of researchers Omar et al. [41] from Malaysia conducted a comparison of visual aids for enhancing reading performance in children with dyslexia. Three specific tools were evaluated: a simple magnifier, which contains a magnifying lens to enlarge text; a visual tracking magnifier, which features a guide line along with a magnifying device that is effective in reducing glare while improving readability; and lastly, a typoscope, which focuses on a line of text by blocking out extra words and glare. A case study involving 80 school children revealed that the visual tracking magnifier (VTM) intervention yielded the most significant improvement after 12 weeks of intervention. Hence, author put forward an idea that a VTM can be used as part of a rehabilitation program for all children with dyslexia.
The authors Liu et al. [42] introduced an structural approach to mitigate traditional method of teaching students with dyslexia. The practice of using the whole word to teach reading and writing for children with dyslexia were proven inefficient. The phonics-based program with special including structured, cumulative, multi-sensorial activities stimulate interest, attention as well as cultivate appropriate skills progressively. Several structural activities were designed to acheive the goal. Ranging activities from introduction of simple vowel, consonant letters to smaller phrases and sentence were introduced. Apart from this use of story-based Fitzroy books, laminated abc cards, a word list containing sequence of vowel, consonant alphabets, multi syllables exercise were introduced. Similarly the physical object or materials were used on this process which includes a set of bottle-caps representing each letter (26) in the alphabet in learning alphabets, tactile mat to practice skywriting, white board and whiteboard markers, exercise book, skipping ropes, hula hopes and many more. 28 total students participated on this experiment which resulted as the experiment had improved children progress in reading, spelling, writing and comprehension significantly determined by pre-test and post-test evaluation.
A group of researcher Ferraz et al. [43] conducted an experiment to determine the effects of a phonological remediation on reading and writing program in children with dyslexia. Researchers evaluated phonological awareness, through multiple exercise like rapid naming, working memory, reading and writing of words and non words, thematic writing. Each phonological intervention consists of exercise involving phonemes and syllables, syllabic and phonemic manipulation, and rhyme and alliteration. The higher complex exercise includes the addition and subtraction of phonemes and syllables, syllabic and phonemic manipulation, oral comprehension etc.
The authors Heikkila and Knight [44] explored inclusive music teaching strategies to enhance the learning abilities of elementary-aged children with developmental dyslexia. By incorporating rhythmic, melodic, and kinesthetic activities, the study aimed to improve phonological awareness, auditory discrimination, and motor coordination. Techniques included using body movements to represent pitches, clapping rhythms, and transitioning these skills to musical instruments like the keyboard. More descriptive categorization of non computing device is illustrated by the Table 8.
Table 8: Descriptive categorization of non-computing devices for children with dyslexia.
Non-computing device Features Benefits
VAKT Strategy Multisensory approach combining visual, auditory, kinesthetic, and tactile modalities. Activities include word tracing, skywriting, writing on the wrist, using word cards, color pens/crayons, mirror work, sand writing, and whiteboard practice. Improves oral reading and reading comprehension by engaging multiple senses.
Visual Tracking Magnifier (VTM) Magnifies text; provides a clear line guide; glare reduction via filters. Enhances reading speed and accuracy, reduces visual strain, and helps maintain line focus.
Simple Magnifier Handheld magnifier that enlarges text without additional visual aids. Improves readability by enlarging text for children with visual tracking challenges.
Typoscope Frames a single line or area of text to reduce visual distractions. Focuses attention on one line at a time, improving comprehension and lowering cognitive load.
SMARTER Phonics Program Structured phonics-based reading program. Materials/activities include story-based Fitzroy readers, laminated ABC cards, syllabified word lists, tactile mat for letter formation, and movement-based tasks (e.g., skipping ropes, hula hoops). Enhances decoding skills, phonemic awareness, and confidence in reading/writing—particularly helpful for rural and indigenous learners.
Phonological Remediation Program Targets phonological awareness, rapid naming, and reading/writing of words and non-words. Exercises include phoneme addition/deletion, syllable segmentation/manipulation, rhyme and alliteration; advanced tasks expand these operations and oral comprehension. Shows significant improvement in reading, writing, and memory retention through systematic intervention.
Music Therapy in Education Integrates melodic and rhythmic activities to build auditory discrimination and engagement. Improves auditory processing, phonological skills, and temporal processing, enhancing overall learning engagement.
[…]
7 Research Questions
7.1 RQ1:What are the current interventions that enhance the learning experience for children with dyslexia, and what are their key features?
The comprehensive literature survey resulted with finding that currently, there are two primary categories of interventions first being technological and second being non-computational. Technological intervention includes application ranging from low tech to high tech devices. These includes, audio books, mobile apps, augmented reality, virtual reality, haptic, tangible user interface, computer apps and natural language processing. Among all of these technological interventions mobile application was found to be most extensively researched domain considering the ease of development and usability among dyslexic children. Most application were focused on creating serious games or game based learning interventions however some application like tangible user interface were focused on providing more realistic way of learning in-cooperating all four (VAKT) multi-sensory parameters. Apart from these, modern immersive technologies like VR and AR systems, have utilized multi sensory approaches to improve letter recognition and phonological skills. Likewise, haptic-based tools have further enhance handwriting fluency and motor coordination through tactile feedback and adaptive guidance.
Non-computing interventions were more common on adopting traditional or prevailing methods. On this method the use of diverse activities such as use of assistive magnifiers, typoscope, play cards, wrist writing, musical games, sand writing etc were implemented that made learning interactive equally important in supporting dyslexic learners. Although classical, it adopts VAKT (Visual, Auditory, Kinesthetic, Tactile) strategy to reinforce memory and improve letter recognition through touch, sound, and movement. On this model structured word exercise were presented which resulted on positive outcome. Furthermore, non-computing intervention presents a challenge of requiring persistent guidance from teacher or caregiver, which makes it more dependent and less efficient comparing the use case of technological devices.
7.2 RQ2:What are the common accessibility features integrated into these interventions for children with dyslexia?
Taking both technological and non-computing devices into account most of the applications targeted the common challenge of dyslexia. This includes enhancing phonological awareness and letter-sound recognition. Immersive technologies like VR and AR were focused on gamified activities that involved object’s and shape identification which were useful training for adhering attention. Non-computing devices provided learning platform, engaging with physical objects like books, colors, cards, whiteboard, it also contains structural exercise to bolster active learning. Different word games like syllabic identification, exercise related to identifying consonant and vowels sounds, spelling test, etc. were common in both technological and non-compositional devices. Many mobile applications were developed, with word games, taking account to the strategies of non-computing methodologies, thereby filling the gap it posses. Many technological devices presented audio cues as assistive features to enhance reading fluency and accuracy, similarly OCR and story based app were common in enhancing readability.
To sum up, gamified learning platform were most common features identified. Integration of rewards, and points to motivate users, real-time feedback were central idea almost every application implemented. Personalized features including customizable application, adjusting system settings such as font size, colors, and word spacing, making reading more comfortable and cross platform support were the next big thing that some application presented. Lastly, every application were implementing VAKT strategy in most possible way to enhance the learning abilities of children with dyslexia.
7.3 RQ3: What gaps exist in current assistive technology research for children with dyslexia, and what opportunities exist for innovation?
Current research on assistive technologies for children with dyslexia posses some gap. First being lack of long-term studies to evaluate the efficacy of this device, there is little to no evidence on whether these benefits persist over time. The existing research were conducted with small number participants, in some case even with their parents. This limits the generalizability of system across diverse population.
On other side, applications integrated with large language model or natural language processing were still limited. Their use case in children with dyslexia might result to a new finding which are yet to explore. Most mobile application were developed which were functional and limited to providing audio and visual cues, however no mobile application had designed an application in-cooperating haptic or vibratory feedback which could further bolster children’s learning, with activated sensory feedback. All this limits the widespread adoption and effectiveness of these technologies.
There are limited applications that offered user personalization, making personalized applications the next domain to explore. Despite these gaps, there are also some opportunities for innovation in this field to consider. First, the integration of haptic feedback into mobile interfaces may enrich user interaction by providing tactile responses, thus improving engagement and accessibility. Second, the incorporation of LLMs and NLP techniques holds potential for improving the comprehension of complex sentences through context-aware simplification and semantic parsing. Lastly, the application of AI in adaptive learning environments presents a compelling direction. AI could enable real-time analysis of a child’s performance and dynamically adjust tasks to tailor their skill level, ensuring personalized and effective learning experiences.
8 Discussion
The systematic literature review highlights the promising role of AR in supporting children with dyslexia. AR applications, which are widely compatible with mobile and tablet devices, are particularly advantageous due to their portability, affordability, and accessibility. These applications often incorporate 3D animated graphics and interactive content, creating immersive, engaging, and multisensory learning environments that cater well to the diverse needs of dyslexic learners
Despite the advancements in assistive technologies, several limitations persist. While tools such as VR and haptic feedback systems demonstrate strong potential for enhancing learning outcomes, their high cost and technical complexity may restrict widespread adoption, especially in low-resource settings. Furthermore, VR systems can contribute to increased screen time, potentially causing eye strain or fatigue, particularly among younger users. As children undergo rapid developmental changes, their learning needs may evolve, which can render static or non-adaptive educational tools less effective over time. These factors underscore the necessity for personalized, adaptable interventions that evolve alongside the child’s cognitive and physiological growth
Similarly, applications that are tailored to learning disabilities possess varying font sizes, and bright background colors which can create a social stigma on use and can discourage them from using assistive technology which further can limit the use case of such devices.The findings from the systematic literature directs researchers or developers on creating more personalized, cross platform technologies in-cooperating multi-sensory modalities for recreating enhancing technologies.
Another significant challenge lies in the design of assistive applications themselves. Tools created specifically for learners with disabilities often feature distinct interface designs—such as enlarged fonts or high-contrast colors—which, while functional, may inadvertently draw attention to the user’s differences. This can contribute to social stigma, ultimately discouraging children from consistently using such tools. Therefore, ensuring that assistive technologies are discreet, inclusive, and socially acceptable is critical for long-term adoption and efficacy. Our work directs researchers or developers on creating more personalized, cross platform technologies in-cooperating multi-sensory modalities for recreating enhancing technologies.
9 Conclusion
The collective findings from the papers on assistive technologies for children with dyslexia highlights the importance of personalized assistive technologies in addressing the unique learning challenges faced by children with dyslexia. Emerging tools such as VR, AR, haptics, and TUI have shown strong potential to enhance motivation and learning outcomes through multisensory (VAKT) engagement. These technologies offer more interactive and tailored experiences. Additionally, non-computing strategies remain effective in improving phonological awareness and memory, serving as valuable alternatives in contexts with limited access to digital resources.
[…]
Acknowledgements
We would like to express sincere gratitude to Professor Krishna Venkatasubramanian of the University of Rhode Island for his valuable guidance guidance, constructive feedback, and continuous support throughout the development of this work.
10 Disclosure of Interest
The authors declare that there are no known competing interests.
11 Disclosure of funding statement
Authors declare that they received no financial aid or funding for this work.
12 Disclosure of Data
This study did not generate any new datasets. All data analyzed in this review were obtained from previously published studies, which are properly cited within the manuscript.